OpenStudy (anonymous):

Jack took a test and got 13 out of 17 points. He took a make-up test and got 12 out of 15 points. The percent of increase in Jack's score is about _____. 3.5% 4.6% 5.9% 6.7%

OpenStudy (anonymous):

3. Q3 Jack took a test and got 13 out of 17 points. He took a make-up test and got 12 out of 15 points. The percent of increase in Jack's score is about _____. Find out the score of each test 13/17 = .7647 = %76.5 12/15 = .8 = 80% 80-76.5=3.5 a. 3.5%